# MiniGPT-5: Interleaved Vision-and-Language Generation via Generative Vokens

### v3

@@ 4.1 Experimental Setup
Baselines  For a comprehensive evaluation of our performance in multimodal generation, we conducted comparative analyses with several prominent baseline models: the Fine-tuned Unimodal Generation Models, Two-stage Baseline, GILL, and Divter.
  • Fine-tuned Unimodal Generation Models: To facilitate fair comparisons in both image and text generation, we fine-tuned two separate models, Stable Diffusion 2.1 and MiniGPT-4 (Zhu et al., 2023), utilizing the VIST dataset. Within the Stable Diffusion 2.1 (Rombach et al., 2022b) model, the U-Net parameters were fine-tuned. For MiniGPT-4’s LLM part, LoRA parameters were fine-tuned.
  • Two-stage Baseline: A common approach in multimodal generation involves first employing Large Language Models (LLMs) to create image captions, which are then fed into text-to-image models for image generation (Wu et al., 2023b). We create such a two-stage baseline for comparison with our end-to-end method by fine-tuning MiniGPT-4 for caption generation and Stable Diffusion 2.1 for text-to-image generation. Given the absence of image descriptions in the VIST dataset, we incorporate a SOTA image captioning model, InstructBLIP-13B (Dai et al., 2023), to generate synthetic captions for supervision.
  • GILL [^1] Given the variations in the valid data within the CC3M dataset, we made adjustments to ensure fair comparisons. Specifically, we retrained it on our specific CC3M data, following the guidelines in their official implementation (https://github.com/kohjingyu/gill). (Koh et al., 2023): GILL is a recent innovation that allows the LLM to generate vokens using a pre-trained text-to-image generation model for single-image generation, where GILL minimizes the Mean Squared Error (MSE) loss between the text-to-image text encoding feature and voken features, similar to $L_{CAP}$ in our approach. For fine-tuning on multimodal datasets, since GILL requires image captions for training, we use Descriptions of Images-in-Isolation (DII) (Huang et al., 2016) in the VIST fine-tuning and generate captions for MMDialog fine-tuning. Contrarily, MiniGPT-5 does not related on all caption data during multimodal generation fine-tuning.
  • Divter (Sun et al., 2021): Divter is a state-of-the-art conversational agent developed for multimodal dialogue contexts. It introduces a customized transformer structure for generating multimodal responses. Divter’s methodology includes pretraining on a vast corpus of text-only dialogues and text-image pairs, followed by fine-tuning on a selected set of multimodal response data. The MMDialog dataset regards Divter’s method as the baseline.
@@
Model CLIP-I ($\uparrow$) FID ($\downarrow$) SD 2.1 (Rombach et al., 2022b) 0.59 393.49 Fine-tuned SD 2.1 0.61 390.25 Two-stage Baseline 0.57 403.06 GILL (Koh et al., 2023) 0.60 381.88 MiniGPT-5 (Prefix Tuning) 0.65 381.55 MiniGPT-5 (LoRA) 0.66 366.62
Model S-BERT ($\uparrow$) Rouge-L ($\uparrow$) Meteor ($\uparrow$) GILL (Koh et al., 2023) 0.3864 0.1784 0.1951 MiniGPT-4 (Zhu et al., 2023) 0.6273 0.3401 0.3296 MiniGPT-5 0.6315 0.3373 0.3263
Metrics  To comprehensively assess the model performance across image, text, and multimodal dimensions, we employ a diverse set of metrics. For evaluating the quality and diversity of generated images, we utilize the Inception Score (IS) (Salimans et al., 2016), and Fréchet Inception Distance (FID) (Heusel et al., 2017). Textual performance is gauged through metrics such as BLEU (Papineni et al., 2002), Rouge-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and Sentence-BERT (S-BERT)  (Reimers & Gurevych, 2019) scores.
From the multimodal perspective, we leverage CLIP-based metrics (Rombach et al., 2022b) to assess the similarities between generated content and ground truth. CLIP-I evaluates the similarity between generated and ground-truth image features. To address potential misalignments in the multimodal generation, such as when the ground truth is text-only, but the output is multimodal, we utilize MM-Relevance (Feng et al., 2022). This metric calculates the F1 score based on CLIP similarities, providing a nuanced evaluation of multimodal coherence.
Recognizing that the generated multimodal output might be meaningful yet differ from the ground truth, we also incorporate human evaluation to assess the model’s performance. We examine the model’s effectiveness from three perspectives: (1) Language Continuity: assessing if the produced text aligns seamlessly with the provided context; (2) Image Quality: evaluating the clarity and relevance of the generated image; and (3) Multimodal Coherence: determining if the combined text-image output is consistent with the initial context.
@@ 4.2 Main Results
In this subsection, we present the performance of different models on the VIST (Huang et al., 2016) and MMDialg (Feng et al., 2022) datasets. Our evaluations span all vision, language, and multimodality domains to showcase the versatility and robustness of the proposed models.
Table 3: Multimodal Story Generation. VIST Human Evaluation on 5,000 samples comparing MiniGPT-5 with both Two-stage Baseline and GILL, across Language Continuity, Image Quality, and Multimodal Coherence aspects. The results highlight the superiority of MiniGPT-5 in more than half cases.
Two-Stage Baseline GILL (Koh et al., 2023)
Model MiniGPT-5 Baseline Tie MiniGPT-5 Baseline Tie
Language Continuity (%) 55.22 34.89 9.89 54.18 35.31 10.51
Image Quality (%) 52.43 37.79 9.78 54.25 35.41 10.34
Multimodal Coherence (%) 56.90 28.88 14.22 55.32 30.34 14.33
Unimodal Generation with Multimodal Input  To evaluate the model performance on image generation and text generation, we systematically provide models with prior history context and subsequently assess the generated images and narrations at each following step. Tables 2 and  2 outline the results of these experiments on the VIST validation set, showing the performance in both image and language metrics, respectively. The findings demonstrate that MiniGPT-5 can generate coherent, high-quality images utilizing long-horizontal multimodal input prompts across all data, without compromising the original model’s ability for multimodal comprehension, indicating the efficacy of our model in diverse settings.
Multimodal Generation with Multimodal Input  To assess the quality of multimodal generation, we test both our model and the baselines on the VIST validation set by human evaluation. Given a preceding multimodal sequence, models are tasked with producing the subsequent scenario for each task. We select a random sample of 5,000 sequences, with each requiring evaluation by two workers. These evaluators are tasked with determining the superior multimodal output based on three criteria: Language Continuity, Image Quality, and Multimodal Coherence. This assessment is facilitated using Amazon Mechanical Turk (Crowston, 2012), with a representative example (Fig. 4) provided in the Appendix. As depicted in Table 3, our model, MiniGPT-5, is found to generate more fitting text narrations in around 55% of instances, deliver superior image quality in around 53% of cases, and produce more coherent multimodal outputs in around 56% of the scenarios. This data distinctly showcases its enhanced multimodal generation capabilities compared to the two-stage baseline, which must generate intermediate image captions first.
@@
Table 4: Multimodal generation results on MMDialog test set. In order to compare with their baseline, we use the same metrics reported in MMDialog (Feng et al., 2022).
Model IS ($\uparrow$) BLEU-1 ($\uparrow$) BLEU-2 ($\uparrow$) Rouge-L ($\uparrow$) MM-Relevance ($\uparrow$)
Divter (Sun et al., 2021) 20.53 0.0944 0.0745 0.1119 0.62
GILL (Koh et al., 2023) 23.78 0.2912 0.1945 0.1207 0.64
@@
Figure 3: Qualitative examples from MiniGPT-5 and baselines on the VIST and MMDialog datasets. The orange blocks indicate the input prompts, while the green blocks include model outputs. The comparisons show that MiniGPT-5 can produce coherent and high-quality multimodal output. We would like to emphasize that MiniGPT-5 does not use any caption data during fine-tuning on VIST and MMDialog, which obeys to our description-free settings. More qualitative examples can be found in the Appendix D.
@@ 4.3 Ablation Studies
To further evaluate the effectiveness of our design, we conducted several ablation studies, and more ablation studies can be found in Appendix C.
Table 5: Evaluation of different method designs for image generation qualities on the CC3M validation set. The results show that all of our designs can help the model better align with the stable diffusion model in the pertaining stage.
@@
Model No Context Text Context Image Context Image-Text Context
SD 2.1 (Rombach et al., 2022b) 0.57 0.59 - -
Fine-tuned SD 2.1 0.59 0.61 - -
Two-stage Baseline 0.54 0.56 0.57 0.58
GILL (Koh et al., 2023) 0.56 0.59 0.60 0.60
MiniGPT-5 (Prefix Tuning) 0.60 0.63 0.68 0.70
MiniGPT-5 (LoRA) 0.61 0.64 0.69 0.70
Text-to-Image Generation Qualities on CC3M
Table 7: Generation Qualities on CC3M and VIST. We find that MiniGPT-5 is better at extracting features from long-horizontal multimodal information than single text input.
CC3M VIST
Stable Diffusion 2.1 (Rombach et al., 2022b) 0.64 26.39 0.59 393.49
[…]
MiniGPT-5 0.61 31.47 0.66 366.62
@@
Instead of multimodal input, we also test single text-to-image generation qualities on the CC3M validation set, as displayed in Table 7. The results indicate that although our model can have better generation on multi-turn multimodal scenarios, Stable Diffusion 2 achieves the best outcomes across all metrics for pure text-to-image generation. Since our model attempts to align with the pretrained text encoder of Stable Diffusion 2 in this stage, there is a slight gap in performance due to the limitation of data amount. Compared with the observations on the VIST dataset, we can conclude that MiniGPT-5 is better at extracting features from long-horizontal multimodal information instead of single text input. This indicates potential future directions on efficiently aligning LLMs with generative models. On the other hand, our model outperforms another state-of-the-art multimodal generation model, GILL, on all metrics, further validating the effectiveness of our design.
5 Conclusion
In this paper, we introduce MiniGPT-5, designed to augment the capabilities of LLMs for multimodal generation by aligning the LLM with a pretrained text-to-image generation model. Our approach demonstrates substantial improvements, as evidenced by comprehensive experiments. There are still some limitations of MiniGPT-5. For example, we still find the object texture hard to maintain in the new generation, and the generated image quality still has space to improve. Through this work, we aspire to set a new benchmark for multimodal generative models, opening doors to applications previously deemed challenging due to the disjointed nature of existing image and text synthesis paradigms.
Appendix A Implementation Details
Appendix B Experimental Settings
B.1 Datasets
CC3M (Sharma et al., 2018): Conceptual Captions (CC3M) dataset represents a remarkable collection of high-quality image captions, amassing approximately 3.3 million pairs of text and images from the internet. The CC3M dataset’s diverse content, quality assurance, and support for multimodal learning make it a valuable asset for researchers and AI enthusiasts. Each dataset sample consists of an image accompanied by a corresponding text description, reflecting the richness of human language and visual perception. However, after accounting for license restrictions and eliminating invalid image links, the dataset comprises approximately 2.2 million data pairs suitable for training purposes and 10 thousand data pairs designated for validation.
VIST (Huang et al., 2016): Visual Storytelling (VIST) dataset is an innovative compilation of visual narratives. The VIST dataset’s engaging content, narrative structure, and emphasis on sequential understanding position it as an essential resource for researchers focusing on sequential image understanding. Each sequence within this dataset consists of five images accompanied by corresponding textual narratives, showcasing the intricate interplay between visual imagery and storytelling. Designed to foster creativity and challenge conventional image-captioning models, the dataset provides a platform for training and validating algorithms capable of generating coherent and contextually relevant stories. After eliminating the invalid image links, we got over 65 thousand unique photos organized into more than 34 thousand storytelling sequences for training and 4 thousand sequences with 8 thousand images for validation.
MMDialog (Feng et al., 2022): Multi-Modal Dialogue (MMDialog) dataset stands as the largest collection of multimodal conversation dialogues. The MMDialog dataset’s extensive scale, real human-human chat content, and emphasis on multimodal open-domain conversations position it as an unparalleled asset for researchers and practitioners in artificial intelligence. Each dialogue within this dataset typically includes 2.59 images, integrated anywhere within the conversation, showcasing the complex interplay between text and visual elements. Designed to mirror real-world conversational dynamics, the dataset is a robust platform for developing, training, and validating algorithms capable of understanding and generating coherent dialogues that seamlessly blend textual and visual information.
B.2 Data Format
Fine-tuning Stage In this stage, we utilize the VIST and MMDialog datasets, which contain multi-turn multimodal data. During training, we integrate placeholders for input images, such as ’$<$Img>>><<<ImageHere>>><<</Img$>$’, into the input text prompts when applicable. These prompts also encompass various instructions corresponding to different task types, with outputs manifesting as pure-text, pure-voken, or text-voken combinations. Below, we present example templates in the VIST dataset to illustrate the different task types:
  • Text Generation: Input: “$<$History Context$>$ What happens in the next scene image: $<$Img>>><<<ImageHere>>><<</Img$>$”; Output: “$<$Text Description$>$
  • Image Generation: Input: “$<$History Context$>$ Generate an image with the scene description: [Text Description]”; Output: “[IMG1]…[IMGn]”
  • Text-Image Generation: Input: “$<$History Context$>$ What should happen then?”; Output: “$<$Text Description$>$ [IMG1]…[IMGn]”
Appendix C More Experiments
C.1 Evaluation of Guidance Scale
(a) FID vs CFG Scale
(b) IS vs CFG Scale
(c) CLIP-T vs CFG Scale
(d) CLIP-I vs CFG Scale
C.2 Evaluation of Voken Number
C.3 Ablation of Model Designs
@@
This section explores alternatives to the transformer encoder/decoder architecture discussed in the main paper. Specifically, we experimented with two additional settings: Fixed Queries, and Decoder-Only model where learnable queries are fed into the transformer decoder. For the fixed queries design, we initialize queries the same as learnable queries experiments in the main paper and keep them fixed during training. In the decoder-only approach, we utilize solely the transformer decoder and apply padding to the decoder’s output, ensuring that the token length reaches 77. This length adjustment allows the output to be compatible with the Stable Diffusion encoder. The results of these experiments are detailed in Table 8. From the results of MiniGPT-5 with fixed queries, we find there exists a slight trade-off between image-text coherence and image qualities, where fixed queries can lead to higher image metrics (IS and FID) but lower CLIP similarities. Meanwhile, MiniGPT-5 consistently outperforms the Decoder-Only results in all four evaluation metrics, validating the robustness and efficacy of MiniGPT-5’s transformer encoder/decoder architecture design.
@@
(a) FID vs $n_{voken}$
(b) IS vs $n_{voken}$
(c) CLIP-T vs $n_{voken}$
(d) CLIP-I vs $n_{voken}$

### v4

@@ 4.2 Experimental Setup
4.2.1 Baselines
For a comprehensive evaluation of our performance in multimodal generation, we conducted comparative analyses with several prominent baseline models: the Fine-tuned Unimodal Generation Models, Two-stage Baseline, GILL  [^1] Given the variations in the valid data within the CC3M dataset, we made adjustments to ensure fair comparisons. Specifically, we retrained it on our specific CC3M data, following the guidelines in their official implementation (https://github.com/kohjingyu/gill). [koh2023generating], and Divter [sun2021multimodal]:
  • Fine-tuned Unimodal Generation Models: To facilitate fair comparisons in both image and text generation, we fine-tuned two separate models, Stable Diffusion 2.1 and MiniGPT-4 [zhu2023minigpt], utilizing the VIST dataset. Within the Stable Diffusion 2.1 [rombach2021highresolution] model, the U-Net parameters were fine-tuned. For MiniGPT-4’s LLM part, LoRA parameters were fine-tuned.
  • Two-stage Baseline: A common approach in multimodal generation involves first employing Large Language Models (LLMs) to create image captions, which are then fed into text-to-image models for image generation [wu2023visual]. We create such a two-stage baseline for comparison with our end-to-end method by fine-tuning MiniGPT-4 for caption generation and Stable Diffusion 2.1 for text-to-image generation. Given the absence of image descriptions in the VIST dataset, we incorporate a SOTA image captioning model, InstructBLIP-13B [instructblip], to generate synthetic captions for supervision.
  • GILL: GILL is a recent innovation that allows the LLM to generate vokens using a pre-trained text-to-image generation model for single-image generation, where GILL minimizes the Mean Squared Error (MSE) loss between the text-to-image text encoding feature and voken features, similar to $L_{CAP}$ in our approach. For fine-tuning on multimodal datasets, since GILL requires image captions for training, we use Descriptions of Images-in-Isolation (DII) [huang2016visual] in the VIST fine-tuning and generate captions for MMDialog fine-tuning. Contrarily, MiniGPT-5 does not related on all caption data during multimodal generation fine-tuning.
  • Divter [sun2021multimodal]: Divter is a state-of-the-art conversational agent developed for multimodal dialogue contexts. It introduces a customized transformer structure for generating multimodal responses. Divter’s methodology includes pretraining on a vast corpus of text-only dialogues and text-image pairs, followed by fine-tuning on a selected set of multimodal response data. The MMDialog dataset regards Divter’s method as the baseline.
Model CLIP-I ($\uparrow$) FID ($\downarrow$) SD 2.1 [rombach2021highresolution] 0.59 393.49 Fine-tuned SD 2.1 0.61 390.25 Two-stage Baseline 0.57 403.06 GILL [koh2023generating] 0.60 381.88 MiniGPT-5 (Prefix Tuning) 0.65 381.55 MiniGPT-5 (LoRA) 0.66 366.62
@@ 4.2.1 Baselines
Model S-BERT ($\uparrow$) Rouge-L ($\uparrow$) Meteor ($\uparrow$) GILL [koh2023generating] 0.3864 0.1784 0.1951 MiniGPT-4 [zhu2023minigpt] 0.6273 0.3401 0.3296 MiniGPT-5 0.6315 0.3373 0.3263
@@
4.2.2 Metrics
To comprehensively assess the model performance across image, text, and multimodal dimensions, we employ a diverse set of metrics. For evaluating the quality and diversity of generated images, we utilize the Inception Score (IS) [salimans2016improved], and Fréchet Inception Distance (FID) [heusel2017gans]. Textual performance is gauged through metrics such as BLEU [papineni2002bleu], Rouge-L [lin2004rouge], METEOR [banerjee2005meteor], and Sentence-BERT (S-BERT)  [reimers2019sentence] scores.
From the multimodal perspective, we leverage CLIP-based metrics [rombach2021highresolution] to assess the similarities between generated content and ground truth. CLIP-I evaluates the similarity between generated and ground-truth image features. To address potential misalignments in the multimodal generation, such as when the ground truth is text-only, but the output is multimodal, we utilize MM-Relevance [feng2022mmdialog]. This metric calculates the F1 score based on CLIP similarities, providing a nuanced evaluation of multimodal coherence.
We also incorporate human evaluation to assess the model’s performance. We examine the model’s effectiveness from three perspectives: (1) Language Continuity: assessing if the produced text aligns seamlessly with the provided context; (2) Image Quality: evaluating the clarity and relevance of the generated image; and (3) Multimodal Coherence: determining if the combined text-image output is consistent with the initial context.
@@ 4.3 Main Results
Figure 3: Qualitative examples from MiniGPT-5 and baselines on the VIST and MMDialog datasets. The orange blocks indicate the input prompts, while the green blocks include model outputs. The comparisons show that MiniGPT-5 can produce coherent and high-quality multimodal output. We would like to emphasize that MiniGPT-5 does not use any caption data during fine-tuning on VIST and MMDialog, which obeys to our description-free settings. More qualitative examples can be found in the Appendix C.
In this subsection, we present the performance of different models on the VIST [huang2016visual] and MMDialg [feng2022mmdialog] datasets. Our evaluations span all vision, language, and multimodality domains to showcase the versatility and robustness of the proposed models.
Model MiniGPT-5 Two-stage Baseline Tie
Language Continuity (%) 55.22 34.89 9.89
Image Quality (%) 52.43 37.79 9.78
Multimodal Coherence (%) 56.90 28.88 14.22
Table 3: VIST Human Evaluation on 5,000 samples for multimodal generation from Language Continuity, Image Quality, and Multimodal Coherence aspects. The results indicate, in more than 70% cases, the MiniGPT-5 is better or on par with the two-stage baseline.
Unimodal Generation on VIST  To evaluate the model performance on image generation and text generation, we systematically provide models with prior history context and subsequently assess the generated images and narrations at each following step. Tables 2 and  2 outline the results of these experiments on the VIST validation set, showing the performance in both image and language metrics, respectively. The findings demonstrate that MiniGPT-5 can generate coherent, high-quality images utilizing long-horizontal multimodal input prompts across all data, without compromising the original model’s ability for multimodal comprehension, indicating the efficacy of our model in diverse settings.
Multimodal Generation on VIST  To assess the quality of multimodal generation, we test both our model and the baselines on the VIST validation set by human evaluation. Given a preceding multimodal sequence, models are tasked with producing the subsequent scenario for each task. We select a random sample of 5,000 sequences, with each requiring evaluation by two workers. These evaluators are tasked with determining the superior multimodal output based on three criteria: Language Continuity, Image Quality, and Multimodal Coherence. This assessment is facilitated using Amazon Mechanical Turk [crowston2012amazon], with a representative example (Fig. 4) provided in the Appendix. As depicted in Table 3, our model, MiniGPT-5, is found to generate more fitting text narrations in around 55% of instances, deliver superior image quality in around 53% of cases, and produce more coherent multimodal outputs in around 56% of the scenarios. This data distinctly showcases its enhanced multimodal generation capabilities compared to the two-stage baseline, which must generate intermediate image captions first.
@@
Model IS ($\uparrow$) BLEU-1 ($\uparrow$) BLEU-2 ($\uparrow$) Rouge-L ($\uparrow$) $M$-Relevance ($\uparrow$)
Divter [sun2021multimodal] 20.53 0.0944 0.0745 0.1119 0.62
GILL [koh2023generating] 23.78 0.2912 0.1945 0.1207 0.64
@@
Table 4: Multimodal generation results on MMDialog test set. In order to compare with their baseline, we use the same metrics reported in MMDialog [feng2022mmdialog].
@@ 4.4 Ablation Studies
To further evaluate the effectiveness of our design, we conducted several ablation studies, and more ablation studies can be found in Appendix B.
@@
Table 5: Evaluation of different method designs for image generation qualities on the CC3M validation set.
@@
Model No Context Text Context Image Context Image-Text Context
SD 2 [rombach2021highresolution] (Zero-shot) 0.57 0.59 - -
GILL [koh2023generating] (Zero-shot) 0.54 0.54 0.55 0.54
MiniGPT-5 (Zero-shot) 0.54 0.57 0.57 0.57
Fine-tuned SD 2 0.59 0.61 - -
Two-stage Baseline 0.54 0.56 0.57 0.58
MiniGPT-5 (Prefix Tuning) 0.60 0.63 0.68 0.70
MiniGPT-5 (LoRA) 0.61 0.64 0.69 0.70
@@
Influence of Model Designs for Image Generation  To validate our feature mapper encoder/decoder architecture, we tested two alternatives: a model with Fixed Queries, where queries are initialized but not trained, and a Decoder-Only model, which pads its output for compatibility with the Stable Diffusion encoder. The results of these experiments are detailed in Table 7. From the results of MiniGPT-5 with fixed queries, we find there exists a slight trade-off between image-text coherence and image qualities, where fixed queries can lead to higher image metrics (IS and FID) but lower CLIP similarities. Meanwhile, MiniGPT-5 consistently outperforms the Decoder-Only results in all four evaluation metrics, validating the robustness and efficacy of MiniGPT-5’s transformer encoder/decoder architecture design.
@@
CC3M VIST
Stable Diffusion 2.1 [rombach2021highresolution] 0.64 26.39 0.59 393.49
Stable Diffusion 3 [sd3] 0.68 29.19 0.63 380.35
MiniGPT-5 (MiniGPT-4 + SD 2.1) 0.61 31.47 0.66 366.62
MiniGPT-5 (LLaVA-1.5 + SD 2.1) 0.62 28.96 0.65 376.58
MiniGPT-5 (Qwen2.5-VL + SD3) 0.64 29.31 0.70 340.20
Table 8: Backbone/base-model comparison on CC3M and VIST. SD 2.1/SD3 are unimodal T2I baselines; MiniGPT-5 rows are interleaved text–image generation.
Backbone and Base-Model Study We evaluate whether the interface of MiniGPT-5 depends on a particular multimodal backbone or image generator. We consider three MLLMs, MiniGPT-4 [zhu2023minigpt], LLaVA-1.5 [liu2024improved], and Qwen2.5-VL [qwen2_5_vl], and two diffusion backbones, Stable Diffusion 2.1 (SD 2.1) [rombach2021highresolution] and Stable Diffusion 3 (SD3) [sd3], shown in Table 8. On VIST, MiniGPT-5 (Qwen2.5-VL+SD3) outperforms both SD3 alone and MiniGPT-5 (MiniGPT-4+SD2.1), indicating that stronger backbones translate into better interleaved coherence through our voken-to-diffusion interface. With the generator fixed (SD 2.1), swapping MiniGPT-4 for LLaVA-1.5 yields comparable VIST results and a modest improvement on CC3M, suggesting robustness to the choice of MLLM. On CC3M, unimodal T2I baselines achieve the best single-turn extremes, while MiniGPT-5 (Qwen2.5-VL+SD3) remains competitive and uniquely supports interleaved generation; the residual gap likely stems from the added mapping from LLM states to diffusion conditions and our objective emphasizing interleaving rather than single-turn T2I. Overall, gains on VIST persist across MLLMs and improve with a stronger generator, supporting that our contributions are orthogonal to specific backbones yet benefit from newer ones.
5 Conclusion
We introduce MiniGPT-5, designed to augment the capabilities of LLMs for multimodal generation by aligning the LLM with a pretrained text-to-image generation model. Our approach demonstrates substantial improvements. The limitation of MiniGPT-5 is that we still find the object texture is hard to maintain in the new generation. Through this work, we aspire to set a new benchmark for existing and future multimodal generative models, opening doors to applications previously deemed challenging due to the disjointed nature of existing image and text synthesis paradigms.
Appendix A Experimental Settings
A.1 Datasets
CC3M [sharma2018conceptual]: Conceptual Captions (CC3M) dataset represents a remarkable collection of high-quality image captions, amassing approximately 3.3 million pairs of text and images from the internet. The CC3M dataset’s diverse content, quality assurance, and support for multimodal learning make it a valuable asset for researchers and AI enthusiasts. Each dataset sample consists of an image accompanied by a corresponding text description, reflecting the richness of human language and visual perception. However, after accounting for license restrictions and eliminating invalid image links, the dataset comprises approximately 2.2 million data pairs suitable for training purposes and 10 thousand data pairs designated for validation.
VIST [huang2016visual]: Visual Storytelling (VIST) dataset is an innovative compilation of visual narratives. The VIST dataset’s engaging content, narrative structure, and emphasis on sequential understanding position it as an essential resource for researchers focusing on sequential image understanding. Each sequence within this dataset consists of five images accompanied by corresponding textual narratives, showcasing the intricate interplay between visual imagery and storytelling. Designed to foster creativity and challenge conventional image-captioning models, the dataset provides a platform for training and validating algorithms capable of generating coherent and contextually relevant stories. After eliminating the invalid image links, we got over 65 thousand unique photos organized into more than 34 thousand storytelling sequences for training and 4 thousand sequences with 8 thousand images for validation.
MMDialog [feng2022mmdialog]: Multi-Modal Dialogue (MMDialog) dataset stands as the largest collection of multimodal conversation dialogues. The MMDialog dataset’s extensive scale, real human-human chat content, and emphasis on multimodal open-domain conversations position it as an unparalleled asset for researchers and practitioners in artificial intelligence. Each dialogue within this dataset typically includes 2.59 images, integrated anywhere within the conversation, showcasing the complex interplay between text and visual elements. Designed to mirror real-world conversational dynamics, the dataset is a robust platform for developing, training, and validating algorithms capable of understanding and generating coherent dialogues that seamlessly blend textual and visual information.
A.2 Data Format
Fine-tuning Stage In this stage, we utilize the VIST and MMDialog datasets, which contain multi-turn multimodal data. During training, we integrate placeholders for input images, such as ’<<Img>><<ImageHere>><</Img>>’, into the input text prompts when applicable. These prompts also encompass various instructions corresponding to different task types, with outputs manifesting as pure-text, pure-voken, or text-voken combinations. Below, we present example templates in the VIST dataset to illustrate the different task types:
  • Text Generation: Input: “<<History Context>> What happens in the next scene image: <<Img>><<ImageHere>><</Img>>”; Output: “<<Text Description>>
  • Image Generation: Input: “<<History Context>> Generate an image with the scene description: [Text Description]”; Output: “[IMG1]…[IMGn]”
  • Text-Image Generation: Input: “<<History Context>> What should happen then?”; Output: “<<Text Description>> [IMG1]…[IMGn]”
Appendix B More Experiments
B.1 Evaluation of Guidance Scale
B.2 Evaluation of Voken Number
B.3 Texture Preservation via FID
To assess whether models preserve the textures present in prior images when generating the last image of a story, we evaluate on VIST in a “final–step” setting: for each story, all preceding step images and narrations are provided as context, and the model must produce the final image. We then compute FID between the set of generated final images and the ground-truth final images from VIST. This directly measures the realism and low-level appearance consistency of the predicted finals relative to the true finals under an identical multimodal context, shown in Table 9. In this setting, ViLGen (Qwen2.5-VL + SD3) attains the lowest FID, indicating stronger retention of low-level textures in the generated final images.
Model FID ($\downarrow$)
GILL [koh2023generating] 61.85
MiniGPT-5 (MiniGPT-4 + SD 2.1) 59.48
MiniGPT-5 (Qwen2.5-VL + SD3) 56.32
Table 9: Texture preservation on VIST (final–step).